OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (1, 9) and (-1, 11)? A. y + 9 = -(x - 1) B. y - 9 = 1(x + 1) C. y + 9 = 1(x + 1) D. y - 9 = -(x - 1)

OpenStudy (anonymous):

first we need to find the slope by using the slope formula : slope(m) = (y2 - y1) / (x2 - x1) (1,9) x1 = 1 and y1 = 9 (-1,11) x2 = -1 and y2 = 11 now we sub slope(m) = (11 - 9) / (-1-1) m = 2/-2 = -1 <== slope is -1 now we will use point slope formula : y - y1 = m(x - x1) (slope)m) = -1 (1,9) x1 = 1 and y1 = 9 now we sub y - 9 = -1(x - 1) or y - 9 = -(x - 1) any questions ?

OpenStudy (jojo4eva):

so it can equal Y= -x+10

OpenStudy (jojo4eva):

lol @kelliegirl33 i was about to hit post when yours came up so i just deleted mine lol

OpenStudy (jojo4eva):

therefore the answer would be A